there are 60 votes in an election
there are two candidates, jerry and bill.
jerry had 18 more votes than bill
how many votes did bill get?

Answer by ikleyn(52858)

.

J + B = 60

J - B = 18

-------------------  Add the equations


2J    = 60 + 18 = 78


J = 78/2 = 39.


Then from the first equation  B = 60-J = 60 - 39 = 21.


ANSWER.  Jerry 39;  Bill 21.

Solved.